Find the sum of: (a) 137 and – 354  
gfjhdnrfnydtjnrfhr

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

(137) + (-354) = (317) + (-137) + (-217)

= 0 + (-217) [(317) + (-317) = 0] = (-217)


Thanks